# OSV: One Step is Enough for High-Quality Image to Video Generation

### v1

@@ Abstract
Video diffusion models have shown great potential in generating high-quality videos, making them an increasingly popular focus. However, their inherent iterative nature leads to substantial computational and time costs. While efforts have been made to accelerate video diffusion by reducing inference steps—through techniques like consistency distillation and GAN training—these approaches often fall short in either performance or training stability. In this work, we introduce a two-stage training framework that effectively combines consistency distillation with GAN training to address these challenges. Additionally, we propose a novel video discriminator design, which eliminates the need for decoding the video latents and improves the final performance. Our model is capable of producing high-quality videos in merely one-step, with the flexibility to perform multi-step refinement for further performance enhancement. Our quantitative evaluation on the OpenWebVid-1M benchmark shows that our model significantly outperforms existing methods. Notably, our 1-step performance (FVD $171.15$) exceeds the 8-step performance of the consistency distillation based method, AnimateLCM (FVD $184.79$), and approaches the 25-step performance of advanced Stable Video Diffusion (FVD $156.94$).
Introduction
@@
Video synthesis provides rich visual effects and creative expression for films, television, advertisements, and games. Diffusion models are playing an increasingly important role in video synthesis (Blattmann et al. 2023c; Shi et al. 2024; Blattmann et al. 2023a; Singer et al. 2022; Ho et al. 2022). Typically, diffusion models involve a forward process and a reverse process. In the forward process, real data is iteratively perturbed with noise until it converges to a simple noise distribution, typically Gaussian. In the reverse process, the noise is gradually removed, ultimately transitioning back to the target data distribution. However, this reverse process usually requires the numerical solution of a generative ODE, termed Probability Flow ODE (PF-ODE) (Song et al. 2020). The iterative nature of this numerical solving process leads to significantly higher computational costs compared to other generative models (e.g., GANs) (Goodfellow et al. 2014). These computational demands are even more significant in video synthesis; for instance, generating a short 2-second video clip using Stable Video Diffusion (SVD) (Blattmann et al. 2023a) on a high-performance A100 GPU can take over 30 seconds.
@@
Figure 1: OSV is a two-stage video diffusion acceleration strategy. In the first stage, we apply GAN for better training efficiency. In the second stage, we apply consistency distillation to boost the performance upper-bound.
@@
Figure 2: Summative motivation. We observe and summarize crucial limitations for (latent) consistency models, and generalize the design space, well tackling all these limitations.
@@
Currently, several strategies have been proposed to reduce the computational costs associated with video generation. Some approaches utilize consistency distillation in the latent space (LCM) (Luo et al. 2023; Song et al. 2023) for acceleration; however, they often struggle to achieve competitive results in few-step settings, such as one or two steps. Other methods initialize with pre-trained diffusion models and incorporate adversarial loss in a GAN-based framework to accelerate generation (GAN). Nonetheless, these methods frequently suffer from training instability. Moreover, since advanced diffusion models are typically trained in the latent space (i.e., they apply an auto-encoder to encode the high-resolution videos into the latent space to facilitate the training), GAN-based methods require decoding latent features into the real image or video space before passing them to the discriminator. This process incurs substantial memory usage and computational overhead, particularly in high-resolution generation tasks. Figure 2 illustrates our summative observations of previous methods, which we summarize as follows:
@@
Exposure Issues: Although LCM can accept classifier-free guidance (CFG) (Ho and Salimans 2021), higher CFG values can lead to exposure problems (Wang et al. 2024a). This complicates the selection of hyperparameters for training.
@@
Increased Training Burden: Because advanced diffusion models are typically trained in the latent space—where high-resolution videos are encoded via an auto-encoder—GAN-based approaches need to decode these latent features back into the real image or video space before they can be evaluated by the discriminator (Sauer et al. 2023b). This decoding step significantly increases memory consumption and computational load, particularly when dealing with high-resolution or long video-length generation tasks.
@@
In this work, we present OSV (One Step Video Generation), allowing for high-quality image-to-video generation in one step while still supporting multi-step refinement. OSV is a two-stage video diffusion acceleration training strategy. Specifically, in the first stage, we incorporate Low-Rank Adaptation (LoRA) (Hu et al. 2021) and fully utilize GANs training, with real data as the true condition for the GANs, greatly accelerating model convergence. In the second stage, we introduce the LCM training function, with data generated by the teacher model serving as the true condition for the GANs, while only fine-tuning specific layers of the network. To further facilitate the training convergence, we replace the commonly applied one-step ODE solver with multi-step solving, ensuring higher distillation accuracy. The second training stage addresses GANs training instability in the later training phase and further improves the performance upper-bound with the knowledge transferred from the teacher video diffusion models. To further improve the training efficiency, we revise the current popular discriminator designs and propose to discard the VAE decoder in the adversarial training. As illustrated in Figure 3, we replace the VAE decoder with a simple up-sampling operator. That is, we directly feed the upsampled video latent into the discriminator whose backbone is pre-trained on the real image/video space (DINOv2 (Oquab et al. 2023)). Our ablation study shows it not only reduces the training cost but also achieves better performance. Our proposed two-stage training not only facilitates early stage training efficiency but also stablizes training and improves the performance upper-bound in the later stage. As shown in Figure 1 verifies our claim.
@@
Related Works
Diffusion Distillation. Denoising process usually has many steps, making them 2-3 orders of magnitude slower than other generative models such as GANs and VAEs. Recent progress on diffusion acceleration has focused on speeding up iteratively time-consuming generation process through distillation (Song et al. 2023; Song and Dhariwal 2023; Salimans and Ho 2022; Heek, Hoogeboom, and Salimans 2024; Yan et al. 2024; Ren et al. 2024; Zheng et al. 2024; Sauer et al. 2023b, 2024; Kim et al. 2023; Zhou et al. 2024; Wang et al. 2024a). Typically, they train a generator to approximate the ordinary differential equation (ODE) sampling trajectory of the teacher model, resulting in fewer sampling steps. Particularly, Progressive Distillation (Salimans and Ho 2022; Meng et al. 2023) trains the student to predict directions pointing to the next flow locations. ADD (Sauer et al. 2023b) leverages an adversarial loss to ensure high-fidelity image generation. SDXL-Lighting (Lin, Wang, and Yang 2024) combines progressive progressive distillation and adversarial distillation, striking a balance between mode coverage and quality.
When it comes to video distillation, Video-LCM (Wang et al. 2023) builds upon existing latent video diffusion and incorporates consistency distillation techniques, achieving high-fidelity and smooth video synthesis with only four sampling steps. Furthermore, Animate-LCM (Wang et al. 2024b) proposes a decoupled consistency learning strategy that separates the distillation of image generation priors and motion generation priors, enhancing visual quality and training efficiency. SF-V (Zhang et al. 2024) follows diffusion-as-GAN paradigms and proposes single-step video generation by leveraging adversarial training on the SVD (Blattmann et al. 2023b) model. In contrast, our empirical observations show that decoupling adversarial latent training and consistency distillation improves training stability and generation quality. Additionally, instead of using UNet itself as the discriminator feature encoder, we use DINOv2 (Oquab et al. 2023) as a feature extractor, which both improves training efficiency and generation quality. Moreover, a multi-step consistency distillation training strategy is also proposed.
Preliminaries
Diffusion Models. Diffusion mdoel (Ho, Jain, and Abbeel 2020; Song et al. 2020) gradually introduces random noise through the diffusion process, transforming the current state ${\mathbf{x}_{0}}$ into a previous state ${\mathbf{x}_{t}}$.
@@
$d{\mathbf{x}}={f}_{t}({\mathbf{x}})\,dt+g_{t}\,d\boldsymbol{w}$ (1)
where $f_{t}(\mathbf{x})=\frac{d\log\alpha_{t}}{d{t}}\mathbf{x}$ and $g_{t}^{2}=\frac{\mathrm{d}\sigma_{t}^{2}}{dt}-2\frac{d\log\alpha_{t}}{dt}\sigma_{t}^{2}$. ${\alpha}_{t}$ is the predefined scale factor and $\boldsymbol{w}_{t}$ denoting the standard Winer process. $\sigma_{t}$ controls the level of noise.
@@
$d\mathbf{x}=\left({f}_{t}(\mathbf{x})-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right)\,dt$ (2)
@@
Consistency Models. Consistency Models (Song et al. 2023) are built upon the PF-ODE in continuous-time diffusion models. Given a PF-ODE that smoothly transforms data into noise, Consistency Models learn to map any point to the initial point of the trajectory at any time step for generative modeling. The formula can be described as:
${f}:\left(\mathbf{x}_{t},t\right)\mapsto\mathbf{x}_{\epsilon},\quad t\in[\epsilon,T].$ (3)
To ensure the boundary conditions hold for any consistent functions, Consistency Models typically employ skip connections. Suppose we have a free-form deep neural network $F_{\theta}$, which can be formulated as:
$f_{\theta}(\mathbf{x},t)=c_{\text{skip}}(t)\mathbf{x}+c_{\text{out}}(t)F_{\theta}(\mathbf{x},t)$ (4)
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions such that $c_{\text{skip}}(\epsilon)=1$, and $c_{\text{out}}(\epsilon)=0$.
For training Consistency Models, the output is enforced to be the same for any pair belonging to the same PF-ODE trajectory, i.e., $f(\mathbf{x}_{t},t)=f(\mathbf{x}_{t^{\prime}},t^{\prime})$ for all $t,t^{\prime}\in[\epsilon,T]$. To maintain training stability, an Exponential Moving Average (EMA) of the target model is used, given by:
$\mathcal{L}_{\text{CD}}^{N}(\theta,\theta^{-};\phi)=\mathbb{E}\left[\lambda(t_{n})d\left(f_{\theta}(\mathbf{x}_{t_{n+1}},t_{n+1}),f_{\theta^{-}}({\mathbf{x}}_{t_{n}}^{\phi},t_{n})\right)\right]$
@@
Figure 3: Comparison of Different Adversarial Training Methods. SF-V requires the encoder of UNet as the feature extraction part of the discriminator. ADD perform adversarial distillation on raw image pixel, which needs to convert latent to image thorough VAE Decoder. Differently, we directly up-sampling the latents, thus requiring less memory and computational resources due to the absence of VAE Decoder.
Figure 4: Overview of OSV. In the first stage, we combine GAN loss and Huber loss (Song and Dhariwal 2023) for better training efficiency. In the second stage, we use consistency distillation loss to boost the performance upper-bound. $b_{sahpe}$, $t_{shape}$, $c_{shape}$, $h_{shape}$ and $w_{shape}$ represent the batch size, number of frames, color channels, height, and width of the input video, respectively.
Method
@@
Network Components. The OSV training process consists of three main components: a student model with weights $\theta$, a EMA model with weights $\theta^{-}$, a pre-trained teacher model with frozen weights $\phi$, and a discriminator with weights $\psi$, as shown in Figure 4. Specifically, the student and teacher models share the same architecture, with the student model initialized from the teacher model. For the discriminator, we adopt the same structure as StyleGAN-T (Sauer et al. 2023a), utilizing DINOv2 (Oquab et al. 2023). We freeze the pretrained weight of DINOv2 and add trainable temporal discriminator heads and spatial discriminator heads for discrimination of the features extracted DINOv2 inspired by SF-V (Zhang et al. 2024). The temporal discriminator heads are composed of 1D convolution blocks. The spatial discriminators are composed of 2D convolution blocks.
@@
$\mathcal{L}_{\text{OSV}}^{\text{d}_{1}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ =$\text{ReLU}(1-D_{\psi}({\mathbf{x}}_{0}))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n})))$
$\mathcal{L}_{\text{OSV}}^{\text{g}_{1}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ = $\lambda^{LGP}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n})))+d(\mathbf{x}_{0},f_{\theta}(\mathbf{x}_{t_{n}}^{\phi},t_{n}))$
where $D_{\psi}$ is the discriminator, $\text{ReLU}(x)=x$ if $x>0$ else $\text{ReLU}(x)=0$ and $\lambda^{LGP}$ is a hyper-parameter. $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c$ (Song and Dhariwal 2023), where $c>0$ is an adjustable constant. It is also worth noting that during this stage, we load the student model with LoRA training and train for a very short period. This ensures rapid image quality improvement at low steps without causing training instability in the student model. Given the mismatch between the pre-training data distribution and the distillation data distribution, using LoRA ensures that the student model retains most of the teacher model’s knowledge while reducing the degradation of image quality at low steps.
Adversarial Consistency Latent Distillation. In the second stage, we introduce the consistency loss. We initialize the student model and EMA model for this stage using the student model from the first stage. Similarly, we initialize the discriminator for the second stage using the discriminator from the first stage. Unlike AnimateLCM, we solve for $\mathbf{x}_{t_{n}}^{\phi}$ using the following equations:
@@
where $c$ represents the condition, which in our case is an image embedding. $m$ controls the number of steps of the ODE solver. We introduce classifier-free guidance distillation similar to (Luo et al. 2023): $\hat{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)={\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi)+w*({\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)-{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi))$, where $c_{zero}$ represents the image embedding set to zero and $w$ controls the strength. We also optimize the adversarial loss between the student-generated data $f_{\theta}(\mathbf{x}_{t_{n+m}})$ and the teacher-generated data $f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})$, as follows:
$\mathcal{L}_{\text{OSV}}^{\text{d}_{2}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ =$\text{ReLU}(1-D_{\psi}(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))\,,$ (9)
$\mathcal{L}_{\text{OSV}}^{\text{g}_{2}}(\boldsymbol{\theta},\boldsymbol{\theta}^{-};\boldsymbol{\phi};\boldsymbol{\psi})$ = $\lambda^{ACD}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))+\lambda(t_{n})d(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n}),f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m}))\,.$ (10)
where $\lambda^{ACD}$ is a hyper-parameter, $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c,c>0$ and $\lambda(t_{n})$ is a weighting function, similar to (Song and Dhariwal 2023). It is worth noting that, as shown in Figure 5, our multi-step solving method not only achieves higher accuracy than the single-step solving method with the same number of iterations but also demonstrates higher accuracy within the same training time. (When $m$ is set to 5, the time for 22,000 iterations is approximately equal to the time for 42,000 iterations when $m$ is set to 1.)

### v2

@@ Abstract
Video diffusion models have shown great potential in generating high-quality videos, making them an increasingly popular focus. However, their inherent iterative nature leads to substantial computational and time costs. Although techniques such as consistency distillation and adversarial training have been employed to accelerate video diffusion by reducing inference steps, these methods often simply transfer the generation approaches from Image diffusion models to video diffusion models. As a result, these methods frequently fall short in terms of both performance and training stability. In this work, we introduce a two-stage training framework that effectively combines consistency distillation with adversarial training to address these challenges. Additionally, we propose a novel video discriminator design, which eliminates the need for decoding the video latents and improves the final performance. Our model is capable of producing high-quality videos in merely one-step, with the flexibility to perform multi-step refinement for further performance enhancement. Our quantitative evaluation on the OpenVid-1M benchmark shows that our model significantly outperforms existing methods. Notably, our 1-step performance (FVD $171.15$) exceeds the 8-step performance of the consistency distillation based method, AnimateLCM (FVD $184.79$), and approaches the 25-step performance of advanced Stable Video Diffusion (FVD $156.94$).
 Equal Contribution
 Project Leader
 Corresponding Author (This work was supported by the Natural Science Foundation of China under contract 62171139).
1 Introduction
@@
Video synthesis provides rich visual effects and creative expression for films, television, advertisements, and games. Diffusion models are playing an increasingly important role in video synthesis [3, 26, 2, 27, 9, 5, 14]. Typically, diffusion models involve a forward process and a reverse process. In the forward process, real data is iteratively perturbed with noise until it converges to a simple noise distribution, typically Gaussian. In the reverse process, the noise is gradually removed, ultimately transitioning back to the target data distribution. However, this reverse process usually requires the numerical solution of a generative ODE, termed Probability Flow ODE (PF-ODE) [29]. The iterative nature of this numerical solving process leads to significantly higher computational costs compared to other generative models (e.g., GANs) [4, 38, 37]. These computational demands are even more significant in video synthesis; for instance, generating a short 2-second video clip using Stable Video Diffusion (SVD) [2] on a high-performance A100 GPU can take over 30 seconds.
@@
Figure 1: OSV is a two-stage video diffusion acceleration strategy. In the first stage, GAN is applied for better training efficiency. In the second stage, we apply consistency distillation to boost the performance upper-bound.
@@
Figure 2: Summative motivation. We observe and summarize crucial limitations for (latent) consistent models and generalize to the design space, which are well addressed by our approach.
@@
Currently, several strategies have been proposed to reduce the computational costs associated with video generation. Some approaches utilize consistency distillation in the latent space (LCM) [16, 30] for acceleration; however, they often struggle to achieve competitive results in few-step settings, such as one or two steps. Other methods initialize with pre-trained diffusion models and incorporate adversarial loss in a GAN-based framework to accelerate generation (GAN). Nonetheless, these methods frequently suffer from training instability. Moreover, since advanced diffusion models are typically trained in the latent space (i.e., they apply an auto-encoder to encode the high-resolution videos into the latent space to facilitate the training), GAN-based methods require decoding latent features into the real image or video space before passing them to the discriminator. This process incurs substantial memory usage and computational overhead, particularly in high-resolution generation tasks. Figure 2 illustrates our summative observations of previous methods, which we summarize as follows:
@@
Exposure Issues: Although LCM can accept classifier-free guidance (CFG) [7], higher CFG values can lead to exposure problems [32]. This complicates the selection of hyperparameters for training.
@@
Increased Training Burden: Because advanced diffusion models are typically trained in the latent space—where high-resolution videos are encoded via an auto-encoder—GAN-based approaches need to decode these latent features back into the real image or video space before they can be evaluated by the discriminator [23]. This decoding step significantly increases memory consumption and computational load, particularly when dealing with high-resolution or long video-length generation tasks.
@@
Moreover, most existing acceleration methods for video diffusion models are derived from image diffusion models (or minor modifications such as adding temporal discriminator heads). For example, Animate-LCM [33] draws inspiration from LCM, and SF-V [39] is influenced by UFOGen [35]. Directly transferring methods such as ADD [23], LADD [24], and UFOGen to video diffusion models presents several issues, including a significant increase in GPU memory usage and the potential for model collapse due to adversarial training, which may lead to extremely weak actions in the generated videos.
In this work, we present OSV (One Step Video Generation), allowing for high-quality image-to-video generation in one step while still supporting multi-step refinement. OSV is a two-stage video diffusion acceleration training strategy. Specifically, in the first stage, we incorporate Low-Rank Adaptation (LoRA) [10] and fully utilize GANs training, with real data as the true condition for the GANs, greatly accelerating model convergence. In the second stage, we introduce the LCM training function, with data generated by the teacher model serving as the true condition for the GANs, while only fine-tuning specific layers of the network. To further facilitate the training convergence, we replace the commonly applied one-step ODE solver with multi-step solving, ensuring higher distillation accuracy. The second training stage addresses GANs training instability in the later training phase and further improves the performance upper-bound with the knowledge transferred from the teacher video diffusion models. To further improve the training efficiency, we revise the current popular discriminator designs and propose to discard the VAE decoder in the adversarial training. As illustrated in Figure 3, we replace the VAE decoder with a simple up-sampling operator. That is, we directly feed the upsampled video latent into the discriminator whose backbone is pre-trained on the real image/video space (DINOv2 [19]). Our ablation study shows it not only reduces the training cost but also achieves better performance. Our proposed two-stage training not only facilitates early stage training efficiency but also stablizes training and improves the performance upper-bound in the later stage. As shown in Figure 1 verifies our claim.
@@
2 Related Works
Diffusion Distillation. Denoising process usually has many steps, making them 2-3 orders of magnitude slower than other generative models such as GANs and VAEs. Recent progress on diffusion acceleration has focused on speeding up iteratively time-consuming generation process through distillation [30, 28, 21, 6, 36, 20, 40, 23, 24, 11, 41, 32, 13]. Typically, they train a generator to approximate the ordinary differential equation (ODE) sampling trajectory of the teacher model, resulting in fewer sampling steps. Particularly, Progressive Distillation [21, 17] trains the student to predict directions pointing to the next flow locations. ADD [23] leverages an adversarial loss to ensure high-fidelity image generation. SDXL-Lighting [15] combines progressive distillation and adversarial distillation, striking a balance between mode coverage and quality.
When it comes to video distillation, Video-LCM [34] builds upon existing latent video diffusion and incorporates consistency distillation techniques, achieving high-fidelity and smooth video synthesis with only four sampling steps. Furthermore, Animate-LCM [33] proposes a decoupled consistency learning strategy that separates the distillation of image generation priors and motion generation priors, enhancing visual quality and training efficiency. SF-V [39] follows diffusion-as-GAN paradigms and proposes single-step video generation by leveraging adversarial training on the SVD [1] model. In contrast, our empirical observations show that decoupling adversarial latent training and consistency distillation improves training stability and generation quality. Additionally, instead of using UNet itself as the discriminator feature encoder, we use DINOv2 [19] as a feature extractor, which both improves training efficiency and generation quality. Moreover, a multi-step consistency distillation training strategy is also proposed.
3 Preliminaries
Diffusion Models. Diffusion mdoel [8, 29] gradually introduces random noise through the diffusion process, transforming the current state ${\mathbf{x}_{0}}$ into a previous state ${\mathbf{x}_{t}}$.
@@ 3 Preliminaries
$d{\mathbf{x}}={f}_{t}({\mathbf{x}})\,dt+g_{t}\,d\bm{w}.$ (1)
where $f_{t}(\mathbf{x})=\frac{d\log\alpha_{t}}{d{t}}\mathbf{x}$ and $g_{t}^{2}=\frac{\mathrm{d}\sigma_{t}^{2}}{dt}-2\frac{d\log\alpha_{t}}{dt}\sigma_{t}^{2}$, ${\alpha}_{t}$ is the predefined scale factor and $\bm{w}_{t}$ denoting the standard Winer process. $\sigma_{t}$ controls the level of noise.
@@
$d\mathbf{x}=\left({f}_{t}(\mathbf{x})-\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right)\,dt.$ (2)
@@
Consistency Models. Consistency Models [30] are built upon the PF-ODE in continuous-time diffusion models. Given a PF-ODE that smoothly transforms data into noise, Consistency Models learn to map any point to the initial point of the trajectory at any time step for generative modeling. The formula can be described as:
${f}:\left(\mathbf{x}_{t},t\right)\mapsto\mathbf{x}_{\kappa},\quad t\in[\kappa,T].$ (3)
$\kappa$ is a number greater than 0 but close to 0. To ensure the boundary conditions hold for any consistent functions, Consistency Models typically employ skip connections. Suppose we have a free-form deep neural network $F_{\theta}$, which can be formulated as:
$f_{\theta}(\mathbf{x},t)=c_{\text{skip}}(t)\mathbf{x}+c_{\text{out}}(t)F_{\theta}(\mathbf{x},t),$ (4)
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions such that $c_{\text{skip}}(\kappa)=1$, and $c_{\text{out}}(\kappa)=0$, $F_{\theta}(\mathbf{x},t)$ represents the output of the neural network.
For training Consistency Models, the output is enforced to be the same for any pair belonging to the same PF-ODE trajectory, i.e., $f(\mathbf{x}_{t},t)=f(\mathbf{x}_{t^{\prime}},t^{\prime})$ for all $t,t^{\prime}\in[\kappa,T]$. To maintain training stability, an Exponential Moving Average (EMA) of the target model is used, given by:
$\mathcal{L}_{\text{CD}}^{N}(\theta,\theta^{-};\phi)=\mathbb{E}\left[\lambda(t_{n})d\left(f_{\theta}(\mathbf{x}_{t_{n+1}},t_{n+1}),f_{\theta^{-}}({\mathbf{x}}_{t_{n}}^{\phi},t_{n})\right)\right],$
@@
Figure 3: Comparison of Different Adversarial Training Methods. SF-V requires the encoder of UNet as the feature extraction part of the discriminator. ADD perform adversarial distillation on raw image pixel, which needs to convert latent to image thorough VAE Decoder. In contrast, we directly upsample the latent signal, replacing the decoder with a simple upsampling layer. Only this modification results in a significant speedup in training on NVIDIA H800 at a resolution of 512$\times$512, reducing the average iteration time from 4.29 seconds to 2.61 seconds, and also decreases the occurrence of floating-point overflows during half-precision training. In addition, OSV training consumes 35.8 gigabytes (GB) of GPU memory, a substantial reduction compared to SF-V’s 73.5 gigabytes (GB) requirement.
Figure 4: Overview of OSV. In the first stage, we combine GAN loss and Huber loss [28] for better training efficiency. In the second stage, we use consistency distillation loss to boost the performance upper-bound. $b_{sahpe}$, $t_{shape}$, $c_{shape}$, $h_{shape}$ and $w_{shape}$ represent the batch size, number of frames, color channels, height, and width of the input video, respectively.
4 Method
@@
Network Components. The OSV training process consists of three main components: a student model with weights $\theta$, a EMA model with weights $\theta^{-}$, a pre-trained teacher model with frozen weights $\phi$, and a discriminator with weights $\psi$, as shown in Figure 4. Specifically, the student and teacher models share the same architecture, with the student model initialized from the teacher model. For the discriminator, we adopt the same structure as StyleGAN-T [22], utilizing DINOv2 [19]. We freeze the pretrained weight of DINOv2 and add trainable temporal discriminator heads and spatial discriminator heads for discrimination of the features extracted DINOv2 inspired by SF-V [39]. The temporal discriminator heads are composed of 1D convolution blocks. The spatial discriminators are composed of 2D convolution blocks.
@@
$\mathcal{L}_{\text{OSV}}^{\text{d}_{1}}(\bm{\theta},\bm{\theta}^{-};\bm{\phi};\bm{\psi})$ =$\text{ReLU}(1-D_{\psi}({\mathbf{x}}_{0}))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n}))),$
$\mathcal{L}_{\text{OSV}}^{\text{g}_{1}}(\bm{\theta},\bm{\theta}^{-};\bm{\phi};\bm{\psi})$ = $\lambda^{LGP}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n}},t_{n})))+d(\mathbf{x}_{0},f_{\theta}(\mathbf{x}_{t_{n}}^{\phi},t_{n})),$
where $D_{\psi}$ is the discriminator, $\text{ReLU}(x)=x$ if $x>0$ else $\text{ReLU}(x)=0$ and $\lambda^{LGP}$ is a hyper-parameter. $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c$ [28], where $c>0$ is an adjustable constant. Notably, during this phase, we load the student model with LoRA training and conduct a very short training period. This approach is adopted because we have observed that full-parameter training tends to result in the generation of static images, which is a consequence of model collapse caused by the input images. Using LoRA ensures that the student model retains most of the knowledge from the teacher model while facilitating rapid convergence. In fact, the first stage can be replaced by loading LoRA modules pretrained with methods such as Animate-LCM, which offers great flexibility. For a fair comparison, we train the first stage using only the method illustrated in Figure 4.
Adversarial Consistency Latent Distillation. LoRA weights stay unmerged in the second stage, part of trainable parameters, and merge only at final inference. We initialize the student model and EMA model for this stage using the student model from the first stage. Similarly, we initialize the discriminator for the second stage using the discriminator from the first stage. Unlike AnimateLCM, we solve for $\mathbf{x}_{t_{n}}^{\phi}$ using the following equations:
@@
where $c$ represents the condition, which in our case is an image embedding. $m$ controls the number of steps of the ODE solver. We introduce classifier-free guidance distillation similar to [16]: $\hat{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)={\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi)+w*({\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c;\phi)-{\Phi}(\mathbf{x}_{t_{n+m}},t_{n+m},c_{zero};\phi))$, where $c_{zero}$ represents the image embedding set to zero and $w$ controls the strength. We also optimize the adversarial loss between the student-generated data $f_{\theta}(\mathbf{x}_{t_{n+m}})$ and the teacher-generated data $f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})$, as follows:
$\mathcal{L}_{\text{OSV}}^{\text{d}_{2}}(\bm{\theta},\bm{\theta}^{-};\bm{\phi};\bm{\psi})$ =$\text{ReLU}(1-D_{\psi}(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n})))+\text{ReLU}(1+D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))\,,$ (9)
$\mathcal{L}_{\text{OSV}}^{\text{g}_{2}}(\bm{\theta},\bm{\theta}^{-};\bm{\phi};\bm{\psi})$ = $\lambda^{ACD}*\text{ReLU}(1-D_{\psi}(f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m})))+\lambda(t_{n})d(f_{\theta^{-}}(\mathbf{x}_{t_{n}}^{\phi},t_{n}),f_{\theta}(\mathbf{x}_{t_{n+m}},t_{n+m}))\,,$ (10)
where $\lambda^{ACD}$ is a hyper-parameter, $d(x,y)=\sqrt{\parallel x-y\parallel^{2}_{2}+c^{2}}-c,c>0$ and $\lambda(t_{n})$ is a weighting function, similar to [28]. It is worth noting that, as shown in Figure 5, our multi-step solving method not only achieves higher accuracy than the single-step solving method with the same number of iterations but also demonstrates higher accuracy within the same training time. (When $m$ is set to 5, the time for 22,000 iterations is approximately equal to the time for 42,000 iterations when $m$ is set to 1.)
@@
Why Decompose into Two Stages? LGP and ACD have fundamental differences. The training objective of LGP is to align the data distribution with the model’s generated distribution. Even when the student model achieves consistency, its adversarial loss remains non-zero, thereby disrupting the consistency learning process. In contrast, ACD does not exhibit this issue. However, LGP matches noise-free data, which facilitates a rapid decrease in the loss during the early stages of training (as shown in Figure 1) and enables the generation of images with less noise in a single step. We follow to the definition of PCM [32], letting $\mathcal{T}_{t\to s}$ denote the flow from $p_{t}$ to $p_{s}$. Let $\mathcal{T}^{\bm{\phi}}_{t\to s}$ and $\mathcal{T}^{\bm{\theta}}_{t\to s}$ represent the transformation mappings of the ODE trajectories for the pre-trained diffusion model and our consistency model, respectively. We redefine the loss functions for LGP and ACD as follows:
$\mathcal{L}_{\text{LGP}}^{adv}(\bm{\theta},\bm{\theta}^{-};\bm{\phi},m)=Dis\left(\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|p_{0}\right)\,,$ (11)
$\mathcal{L}_{\text{ACD}}^{adv}(\bm{\theta},\bm{\theta}^{-};\bm{\phi},m)=Dis\left(\mathcal{T}^{\bm{\theta}}_{t_{n+m}\to\epsilon}\#p_{t_{n+m}}\middle\|\mathcal{T}^{\bm{\theta}^{-}}_{t_{n}\to\epsilon}\mathcal{T}^{\bm{\phi}}_{t_{n+m}\to t_{n}}\#p_{t_{n+m}}\right)\,,$ (12)
where ##\## is the pushforward operator, and $Dis$ is the distribution distance metric. $\mathcal{L}_{\text{LGP}}^{adv}$ is always non-zero, while $\mathcal{L}_{\text{ACD}}^{adv}$ will also converge to zero. We provide a detailed discussion in the appendix.